A circular coil consists 70 closely wound turns and has a radius of 10cm. An externally produced magnetic field of magnitude 2×10^(-3)T is applied perpendicular to the coil.The net flux through the coil is found to vanish when the current in the coil is 2.2A. The inductance of the coil is
Ans 3mH
1
Expert's answer
2020-11-02T09:22:47-0500

Φ = BA

The flux over the coil, equals the flux over one turn multiplied by the number of turns:

Φc = NΦ

Φc = NBA

The area of the coil is given by A = πR2, hence

Φc = πNBR2

"Φc = π(70.0)(2.0 \times 10^{-3}\;T)(10 \times 10^{-2} \;m)^2 = 4.4 \times 10^{-3} \;Wb"

The net magnetic flux through the coil is zero, so the magnetic flux produced by the coil is given by equation:

"L = \frac{Φ_c}{I}"

"L = \frac{4.4 \times 10^{-3} \;Wb}{2.2 \;A} = 2 \times 10^{-3} \;H"

Answer: 2 mH